# Can Large Language Models Really Recognize Your Name?

### v1

@@ 5.1 PII Leakage Detection
We test a total of 10 LLMs, including reasoning (e.g., DeepSeek R1), non-reasoning (e.g., GPT-4o), and small LLMs (e.g., Llama 3.1 8B). To minimize variability, we use a temperature of 0.0 for all of these LLMs whenever applicable. We also use Flair’s large four-class NER model (Akbik et al., 2019), which represents more traditional entity tagging solutions, and a tool called PrivateAI, which represents commercial data leakage detection products.
@@
Moreover, we conduct a small (IRB-exempted) survey in which we asked human volunteers to classify named entities in the 25 templates above. We focus on ambiguous names that LLMs often misclassify, such as Canad, Versache, and Beggiato. We also include the baseline human names and well-known non-human entities in the target ambiguity types as “control” samples, which make up 40% of the survey. We receive 22 responses from members of our institutions. For more details on the survey, see Appendix D.2.
5.2 Abstractive Summarization
We target the Clio system, which is an internal tool at Anthropic for surfacing privacy-preserving insights in Claude users’ conversations (Tamkin et al., 2024). Specifically, we test its conversation summarization and privacy audit module. For this application, we only use the Claude model family (Anthropic, 2024) since Clio is designed around Claude. We evaluate this application using two different types of simulated conversations:
  • Without BPI: The user conversation to be summarized involves asking the model to paraphrase our ambiguous benchmark data (Section 5.1).
  • With BPI: In addition to the normal paraphrasing instructions, we include an instruction at the end to ask the model to “keep the term ‘[NAME]’ intact, even if it looks like a typo.” This combination is designed to overcome Clio’s summarization prompt, which explicitly asks Claude not to include any proper noun.
After running these inputs through the summarization prompt, we then evaluate Clio’s privacy auditor on summaries where the human names are leaked to measure any changes in the auditor’s perceived privacy. We are interested in the following metrics:
  • Summarization leakage %: The percentage of instances where the ambiguous human name is included in the summary. We use McNemar’s test to assess the statistical significance of the changes in leakage before and after BPI.
  • Privacy audit score: The privacy auditor returns an integer score between 1 (“identifiable to an individual”) and 5 (“not identifiable”). We use Wilcoxon Signed-Rank test to assess the statistical significance of the changes in audit scores due to BPI.
6 Results
@@
Sonnet 3.7 0.97 0.00 0.60 0.00 0.81 0.02 0.21 0.00 0.79 0.00 0.68 0.01 0.971 0.00
Instruct DeepSeek V3 0.98 0.08 0.34 0.07 0.87 0.08 0.15 0.00 0.59 0.00 0.59 0.05 0.962 0.00
@@
Small LLMs Gemini 2 Flash 0.98 0.16 0.68 0.22 0.90 0.09 0.39 0.00 0.79 0.00 0.75 0.09 0.995 0.00
@@
Table 1: Recall (R) and false detection rate (FDR) for different ambiguous human name types from our AmBench benchmark. The average is taken over the ambiguous name types. Takeaway: All methods lose $\approx$ 0.2–0.4 points in recall on average for ambiguous human names compared to the baseline.
Here, we describe and analyze the results of the two experiments described in the previous section. Overall, our methods have significant negative impacts on the performance of the PII detection and abstractive summarization applications tested.
6.1 PII Detection
LLMs are much worse at detecting ambiguous human names than popular ones. While most methods achieve nearly perfect recall on the baseline, none of them exceeds 0.8 average recall across the 5 different ambiguous name types and can lose up to 0.4 points as in the case of GPT-4o (Table 1). The best-performing LLM is o3-mini with 0.801 average recall, followed closely by DeepSeek R1 and Llama 3.1 8B. One major reason for the LLMs’ reduced recall is that they confuse the ambiguous human names with the targeted entity types, thus leading to a misclassification or a complete miss (Table 6). For instance, organization-like names tend to be classified as organization, while mineral-like names tend to be not included in the output.
LLMs are not consistent in their detections of ambiguous names. The same name can be assigned different categories depending on the template in which it appears, even though we design the templates to share the same structure and theme. For example, DeepSeek R1, one of the best-performing models on our benchmark, gives inconsistent labels to at least 10% of the names in each ambiguous name type (except for location-like names), particularly so for organization- and mineral-like names (Figure 3). Another example is GPT-4o, which does not consistently classify at least 40% of the names. Whereas for the baseline names, most methods are consistent for $\approx$90% of them.
@@ 6.1 PII Detection
Figure 3: Histograms of the consistency of human name detection for four representative methods (GPT-4o, DeepSeek R1, Llama 3.1 8B, and Flair). Each subfigure corresponds to a different human name type and plots the distribution of the ratio of human name classification for each name across the five templates. Takeaway: Most methods are inconsistent for at least 10% of names in all name types except for the baseline.
Small LLMs have better performance than strong ones, but at the expense of FDR. Bigger non-reasoning LLMs like GPT-4o and DeepSeek V3 have less than 0.6 recall, while LLMs like Qwen 2.5 7B and Gemma 2 9B have at least 0.64 (Table 1). However, the small LLMs’ FDR tends to be at least one or two orders of magnitude larger than the bigger LLMs, particularly with Llama 3.1 8B having nearly 4% FDR on not only ambiguous names but also the baseline. A closer inspection of Llama’s outputs reveals that the model often hallucinates certain normal words as names.
Our human volunteers tend to rely more on the templates than the names, resulting in mixed classification performance (Table 2). While the respondents classify as human names roughly half of the ambiguous name instances (that LLMs tend to misclassify), they do not label 20% of the popular human control samples as humans, but do so for 40% of the non-human control samples. According to our post-survey follow-up questionnaires, nearly two-thirds of the respondents rely more on the templates than the names, and >>>80% are at least moderately influenced by the similarities between the templates.
@@
Ambiguous 0.89 0.53 0.46 0.35 0.49 0.54
Control (human) 0.96 0.96 0.64 0.55 0.82 0.79
Control (non-human) 0.86 0.27 0.23 0.32 0.32 0.40
Table 2: Human volunteers’ recall (i.e., percentage of human name classifications) from our human evaluation study (n = 22). The “Ambiguous” name type refers to AmBench’s ambiguous human names that are often misclassified by LLMs. “Control (human)” refers to popular human names (same as the LLMs’ baseline), while “Control (non-human)” refers to actual non-human entities. Takeaway: Humans can detect some of the personal names that LLMs misclassify but can also miss the baseline human names, which LLMs excel at.
6.2 Abstractive Summarization
[…]

### v2

@@ 5. Evaluating LLMs on AmBench
Experimental Setup   Using AmBench, we test a total of 12 state-of-the-art LLMs, including reasoning (e.g., DeepSeek R1), instruction (e.g., GPT-4o), and small open-weight models (SLM) (e.g., Llama 3.1 8B). To minimize variability, we use a temperature of 0.0 for all of these LLMs whenever applicable. We use the PII detection prompt from the Rescriber system (Zhou et al., 2025) (see Appendix C.2 for the full prompt), which is a complete framework for assisting chatbot users with protecting their prompts by sanitizing PII. Aside from LLMs, we also test Flair’s large four-class NER model (Akbik et al., 2019), which represents more traditional entity tagging solutions, and a tool called PrivateAI [^5] https://private-ai.com/en/redact/, which represents commercial data leakage detection products.
We measure the following metrics:
  • Recall: The ratio of ambiguous human names correctly detected as humans (higher is better). This is measured by checking if the name is contained in the concatenation of all of the LLM’s predictions.
  • Consistency: The percentage of templates where a name is classified as human (higher is better). We say a model is inconsistent with a name if it fails to classify the name as human for any template. Consistency is calculated for each name, and since we only have 5 templates each, the possible values are in $\{0.0,0.2,0.4,0.6,0.8,1.0\}$
  • False Discovery Rate (FDR): The ratio of human name predictions that do not actually match the real human names (lower is better). This is equivalent to 1 - Precision. A high FDR indicates increased hallucination of human names.
@@
Additionally, we conducted a small (IRB-exempted) survey in which we asked human volunteers to classify the names in our text snippets. We focus on ambiguous names that LLMs often misclassify, such as Canad, Versache, and Beggiato. We also include the baseline human names (e.g., Alice) and well-known non-human entities belonging to each target ambiguity source as control samples. For more details on the survey, see Appendix D.
Results   The performance of all 12 LLMs degrades significantly when evaluated on AmBench. More specifically:
Table 1. Recall (R) and False Discovery Rate (FDR) (formatted as percentage) of various LLMs on our AmBench benchmark. The ‘Average’ column is the average of the metrics across the ambiguous name types. The results of the best performing methods in a column are bolded. Takeaway: All methods fail to recognize $\approx 20$4040% on average across the ambiguous human name types.
@@
Gemini 2.5 Pro 0.97 0.03 0.68 0.15 0.87 0.06 0.47 0.00 0.89 0.07 0.78 0.06 0.993 0.00
GPT-5 0.87 0.00 0.59 0.04 0.82 0.02 0.12 0.00 0.71 0.04 0.62 0.02 0.974 0.00
Instruct Gemini 2.5 Flash 0.94 0.28 0.59 0.12 0.90 0.23 0.33 0.11 0.78 0.11 0.71 0.17 0.987 0.02
DeepSeek V3 0.98 0.08 0.34 0.07 0.87 0.08 0.15 0.00 0.59 0.00 0.59 0.05 0.962 0.00
@@
SLMs gpt-oss-20B 0.96 0.48 0.65 0.39 0.96 0.65 0.38 0.73 0.92 0.42 0.77 0.53 0.991 0.00
Gemma 2 9B 0.97 0.23 0.75 0.56 0.90 0.32 0.19 0.00 0.59 0.84 0.68 0.39 0.970 0.00
@@
  1. (1)
    LLMs are much worse at detecting ambiguous human names than popular baseline names. While most methods achieve nearly perfect recall on the baseline, almost none of them achieve higher than 0.8 average recall across the 5 different ambiguous name types and can lose $\approx$0.4 points, as in the case of GPT-5 and 4o (Table 1). The best-performing LLM is GPT-5-mini with 0.82 average recall and is followed closely by Gemini 2.5 Pro and DeepSeek R1, which are all reasoning models. Looking into the reasoning trace, we find that the main reason for the LLMs’ poor performance is that they confuse the ambiguous human names with the targeted entity types, thus leading to a misclassification or a complete miss (Table 8). For instance, organization-like names tend to be classified as organization, while mineral-like names are often not even included in the models’ predictions. See Appendix D.3 for a more comprehensive error analysis, including a detailed error breakdown for each LLM.
  2. (2)
    LLMs are inconsistent in their detections of ambiguous names. The same name can be assigned to different categories depending on the template in which it appears, even though the templates share the same structure and theme (Figure 6). For example, GPT-5-mini, the top-performing instruct model, inconsistently labels at least 10% of the names in each category (except for location-like). With baseline human names, most methods are consistent for at least $\approx$90%.
  3. (3)
    Small LLMs have competitive recall but at the expense of FDR. Bigger instruct LLMs like GPT-4o and DeepSeek V3 often have less than 0.6 recall (the only exception is Gemini 2.5 Flash), while small LLMs have at least 0.64 (Table 1), with gpt-oss-20B leading the group at 0.77 recall. However, their FDR tends to be at least one or two orders of magnitude larger than the FDR of the bigger LLMs, particularly with Llama 3.1 8B having nearly 4% FDR on not only ambiguous names but also the baseline. A closer inspection of Llama’s outputs reveals that the model often hallucinates names.
  4. (4)
    Untrained human volunteers are also impacted by ambiguity (Table 2). While the survey respondents correctly classify nearly 100% of the popular control human names, this figure drops to roughly two-thirds for the ambiguous name instances. 40% of the non-human control samples were tagged as humans, particularly for location- and syndrome-like templates, likely due to the use of the “she” pronoun in location templates and the fact that syndromes are named after humans.
  5. (5)
    Name popularity can be positively correlated with recall, but this relationship varies significantly between the ambiguity sources. For organization-like and bacteria-like names, these two quantities generally show a strong positive correlation (Figure 7). However, for the remaining categories, such positive correlations are only visible for a subset of the LLMs. In fact, at the highest frequency bin, recall can actually decrease. For instance, mineral-like name recall for most LLMs only has a slightly positive correlation with frequency up to near or slightly past the peak of the distribution, after which the recall starts to decrease (like an inverted U shape).
@@
Figure 6. Histograms of the consistency of human name detection for four representative models (GPT-5-mini, Gemini 2.5 Flash, gpt-oss-20B, and Flair). Each sub-figure corresponds to a different ambiguous type, with the x-axis representing the ratio of templates in which a name is classified as human, and the y-axis representing the percentage of names in each category. Takeaway: Most models are inconsistent for at least 10% of AmBench names.
Table 2. Percentage of human name classifications by human survey respondents (n = 204). Takeaway: Humans can detect some of the personal names that LLMs misclassify, but can also miss the baseline human names, which LLMs excel at.
@@
Ambiguous 0.87 0.53 0.79 0.33 0.77 0.66
Control (human) 1.00 0.96 0.88 1.00 1.00 0.97
Control (non-human) 0.71 0.27 0.59 0.23 0.20 0.40
(a) Location
(b) Organization
(c) Syndrome
(d) Bacteria
(e) Mineral
Figure 7. Average recall (y-axis) vs log-transformed number of Google Search results of names (x-axis) for each leakage detection method. Syndrome-like name frequency is based on the first name only. Takeaway: LLMs’ performance is either positively correlated with name frequency for certain orthographic neighbor categories (e.g., organization and bacteria) or has an inverted U-shaped relationship with the frequency.
6. How Well Do Enterprise-grade Solutions Perform? A Case Study with Anthropic’s Clio
Background and Motivation   Clio (Tamkin et al., 2024) (Claude Insights and Observations) is an LLM-powered analytical platform developed by Anthropic AI to characterize how their Claude language model is used at scale by tens of millions of users. The system aims to surface privacy-preserving insights into Claude usage patterns (e.g., topic, language, interaction type) without involving manual human inspection of the raw conversation logs. Clio’s privacy-focused data preprocessing pipeline can be summarized in four main steps (more details can be found in Section 2.3 of Clio’s original paper (Tamkin et al., 2024)):
  1. (1)
    Conversation summarization: Clio prompts their Claude model to generate a summary for each selected conversation, with the following instruction to omit private information: “do not include any [PII] like names, locations, phone numbers, email addresses, and […] any proper nouns.” (Appendix C.3.1).
  2. (2)
    Conversation summary clustering: Using sentence embeddings and k-means, Clio organizes all generated summaries from the previous step into clusters, then discards those whose size does not meet a minimum threshold to avoid overly specific clusters.
  3. (3)
    Cluster summarization: Clio prompts Claude to generate a summary for each cluster using a contrastive approach (e.g., pairing a cluster with nearby summaries that do not belong to the cluster). However, unlike step 1, the model is not instructed to omit private information based on our inspection of the exact prompt used, despite the original paper’s claims (Appendix C.3.2).
  4. (4)
    Cluster summary privacy audit: Clio prompts Claude to rate each cluster summary’s privacy on a scale of 1 (“identifiable to an individual”) to 5 (“not identifiable”), then filters out clusters whose privacy score is below 3 (Appendix C.3.3).
Unlike in the previous experiment, which focuses on name recognition, the role of LLMs in this scenario is to perform name redaction. Given Clio’s large operational scale and tight LLM integration, we want to assess how robustly this real-world industry solution can preserve privacy for (hypothetical) conversations that contain AmBench names. Notably, all of Clio’s LLM prompts are publicly available, allowing external researchers to reproduce their results.
Experimental Setup   We evaluate Clio’s conversation summarization and privacy audit modules as they both depend on LLMs for their privacy requirements (the other components of Clio either do not use LLMs or involve privacy.) We only use the Claude models (Anthropic, 2024) since Clio is designed and evaluated with Claude at its core. We provide Clio with two different types of simulated conversations:
  • Without BPI: The user conversation to be summarized involves asking the model to paraphrase our AmBench test cases, e.g., “Help me paraphrase the following text.”
  • With BPI: In addition to the normal paraphrasing instructions, we include a BPI instruction at the end to ask the model to “[m]ake sure to keep the term ¡NAME¿ intact, even if it looks like a typo.” This combination is designed to test if Clio can still follow its original summarization instruction to exclude all proper nouns.
After running these inputs through the summarization prompt, we then evaluate Clio’s privacy auditor on summaries where the human names are leaked to measure any changes in the auditor’s perceived privacy. We measure the following metrics:
  • Summarization leakage %: The percentage of summaries where the ambiguous human name is included. We use McNemar’s test to assess the statistical significance of the changes in leakage before and after BPI.
  • Clio’s privacy audit score: Clio’s privacy auditor returns an integer score between 1 (“identifiable to an individual”) and 5 (“not identifiable”). We use the Wilcoxon Signed-Rank test to assess the statistical significance of the changes in audit scores due to BPI.
Table 3. Performance of Clio’s summarizer and privacy auditor before and after benign prompt injection (BPI). Note that the privacy audit scores (right side) are only calculated for samples where BPI causes a name leakage. The p-values are calculated for the changes in leakage rate (McNemar) and audit score (Wilcoxon Signed-Rank). The % $\geq$ score column refers to the percentage of privacy audit scores that stay the same or increase after BPI. The average is taken over the ambiguous name types. Takeaway: BPI significantly increases the ambiguous human names leakage rate in Clio’s summaries and privacy audits.
@@ 6. How Well Do Enterprise-grade Solutions Perform? A Case Study with Anthropic’s Clio
Location 7.24 29.75 +17.51 $\ll 0.001$ 4.52 1.55 -2.97 13.69 $\ll 0.001$
Organization 0.00 0.81 +0.81 $\ll 0.001$ 5.0 2.28 -2.72 5.30 $\ll 0.001$
Syndrome 0.03 0.09 +0.06 $\ll$ 0.007 5.0 2.57 -2.43 17.39 $\ll 0.001$
Mineral 11.31 35.42 +24.11 $\ll 0.001$ 5.0 3.35 -1.65 12.57 $\ll 0.001$
Bacteria 1.03 24.93 +23.90 $\ll 0.001$ 5.0 1.67 -3.33 2.30 $\ll 0.001$
Average 3.92 18.20 +14.28 N/A 4.90 2.28 -2.62 10.25 N/A
Baseline 7.60 7.58 -0.02 0.971 5.0 1.36 -3.64 0.56 $\ll 0.001$
ResultsTable 3 shows the impact of combining NRB with BPI on Clio’s privacy protection performance. Specifically:
  1. (1)
@@
7. Potential Mitigations
As LLMs become better at reasoning, their performance on AmBench will likely also improve (Figure 2). However, although better reasoning may help LLMs with ambiguous contexts (Yi et al., 2025), this ability comes with significantly increased inference costs. Here, we explore a few cost-saving approaches by modifying the detection prompt with direct hints to roughly approximate the upper bound on the performance of three representative LLMs (GPT-5 mini, Gemini 2.5 Flash, and gpt-oss 20B). Specifically, we inject our knowledge of what the benchmark data looks like into the prompt for the LLMs:
  • Prompt 1 (P1): We add a special instruction to bias the LLM towards name/username detection:
    If a phrase in the text is unusual, stylized, or formatted in a way that could be a NAME or USERNAME, you must carefully consider this possibility.
  • Prompt 2 (P2): In addition to P1, we add a custom “cheat” example at the end:
    Here is an example:
    Input: "I recently visited Italys. She was as charming as always."
    Output: {"results": [{"entity_type": "NAME", "text": "Italys"}]},
    Explanation: "Italys" may look similar to a GEOLOCATION, but there are actual people bearing this exact name.
These two mitigations are somewhat impractical because they describe a vulnerability that would be unknown at test time. Prompt 2 even uses an example that is structurally very similar to the benchmark. Despite the biased instruction to focus on unusual names/usernames and the example, the best LLMs still leave a gap of $\approx$0.05 average recall between the ambiguous and baseline human names (Table 4). Nevertheless, the improved results suggest that integrating LLMs with a knowledge base of known weaknesses or even rare name examples can potentially support LLMs in their privacy assessment (e.g., retrieval-augmented generation) (Fan et al., 2024).
Table 4. Average Recall (R) and False Discovery Rate (FDR) (formatted in percentage) for three representative LLMs with different prompting methods. Takeaway: Recall improves significantly with the cheat example, but remains 5 percentage points below the baseline.
Prompt Method Location Org. Syndrome Mineral Bacteria Average Baseline
@@ 7. Potential Mitigations
GPT-5 mini Orig. 0.98 0.08 0.65 0.06 0.97 0.02 0.58 0.00 0.93 0.00 0.82 0.03 0.996 0.00
P1 0.98 0.09 0.70 0.06 0.97 0.04 0.71 0.06 0.95 0.06 0.86 0.06 0.997 0.00
P2 0.99 0.07 0.79 0.07 0.98 0.03 0.93 0.00 0.98 0.12 0.93 0.06 0.999 0.00
Gemini 2.5 Flash Orig. 0.94 0.28 0.59 0.12 0.90 0.23 0.33 0.11 0.78 0.11 0.71 0.17 0.987 0.02
P1 0.94 0.49 0.69 0.05 0.93 0.17 0.33 0.00 0.78 0.08 0.73 0.16 0.981 0.00
P2 0.99 0.20 0.91 0.12 0.98 0.20 0.83 0.04 0.97 0.06 0.94 0.12 0.996 0.00
gpt-oss 20B Orig. 0.96 0.48 0.65 0.39 0.96 0.65 0.38 0.73 0.92 0.42 0.77 0.53 0.991 0.00
P1 0.95 0.49 0.64 0.37 0.96 0.51 0.43 0.32 0.92 0.35 0.78 0.41 0.992 0.00
P2 0.95 0.44 0.68 0.47 0.97 0.52 0.67 0.41 0.95 0.28 0.84 0.42 0.993 0.00
8. Discussion and Conclusion
Implications for Privacy   Our experimental findings demonstrate the hidden perils of relying on LLMs to build privacy-focused solutions without fully understanding their failure modes. While LLMs’ imperfect privacy reasoning is not an unknown issue (Mireshghallah et al., 2024b), the fact that they can systematically fail at the very first step of recognizing sensitive data has major consequences for downstream dependencies. Any LLM-based privacy mechanism, like automated PII redaction, may inadvertently expose sensitive data, leading to non-compliance with regulatory requirements such as GDPR or CCPA and exposing organizations to legal and financial risks (Consumer Reports, 2017). Moreover, malicious actors can exploit these vulnerabilities to engineer novel attacks to compromise user privacy. To demonstrate this, we sketch a hypothetical attack on Anthropic’s Clio by leveraging NRB and BPI:
  • Assume an attacker $\mathcal{A}$ who has access to Clio’s final outputs (e.g., an Anthropic analyst) and wants to find more information about a Claude user with an ambiguous name that is known to the attacker. Assume the user’s conversation is included in Clio’s inputs, and their name is accidentally leaked in the conversation summary generated by Clio due to NRB and BPI.
  • $\mathcal{A}$, via Claude’s interface or API, creates a large volume of Claude conversations containing the target user’s name and instruction-like texts. These conversations are designed to get the targeted user’s true conversation clustered in the same group as the attacker’s conversations. If this adversarial cluster is successfully created (e.g., meets the minimum cluster size), then Clio’s LLM-powered cluster summarization module will likely include the user’s name in the cluster description.
  • $\mathcal{A}$ searches in the final outputs of Clio for clusters whose description contains the user’s name.
With this setup, a malicious actor can extract private information from users whose name can be confused with another non-sensitive entity. If we cannot assume that the target user is included in Clio, then this attack can be repurposed into a form of membership inference attack (Shokri et al., 2017) to probe the target’s existence in Anthropic’s data. While we assess the likelihood of having all the conditions required for this exploit to succeed as unlikely, its mere possibility means the system’s privacy promises are not tenable. The example thus highlights how LLMs’ imperfect PII detection ability can open up novel attack avenues in systems that rely solely on this technology without incorporating techniques with formal privacy guarantees (e.g., differential privacy (Liu et al., 2025)).
Implications for Fairness   The observed impact of NRB on names that fall outside the dominant linguistic distributions suggests a risk of uneven privacy protection across populations. Specifically, individuals with morphologically ambiguous names due to the orthographic neighborhood effect may be more likely to have their identities exposed. While this property is not a typical demographic attribution, it nonetheless raises fairness concerns that extend beyond privacy applications. When name recognition serves as an upstream component in tasks such as identity verification, content moderation, or hiring pipelines, these failures may propagate and amplify inequities, effectively excluding certain populations from reliable system performance (Sweeney, 2013; Noble, 2018; Consumer Reports, 2017). Moreover, because such vulnerabilities are often opaque to end users and developers alike, affected individuals may have limited recourse or awareness of the source of harm. Our findings thus highlight the need to treat name handling also as a fairness-critical component in AI-powered privacy systems and beyond.
Table 5. Recall and FDR on Reddit usernames for three representative top-performing LLMs.
Name type Ambiguous Baseline